# Representation noising effectively prevents harmful fine-tuning on LLMs

### v1

@@ 5 Mechanistic Analysis
Figure 5: Harmful probe accuracy on (a) base model and attacked model, (b) base model and models trained with RepNoise ($beta=4$) and adversarial Loss, and (c) base model, RepNoise model and an attacked RepNoise model
@@ 6 Related Work
Few works have attempted to defend against HFAs. Meta-learning approaches have been used to reduce the fine-tunability of Language Models for harmful classification tasks [40] and prevent image classification models from being fine-tuned on restricted domains [41]. However, meta-learning approaches can be uninterpretable and too computationally expensive for LLMs. [28] added a security vector during training to trick a model into thinking that it has already learned the harmful task. [24] keep embeddings close to the original embeddings by adding a perturbation loss called ‘Vaccination’. While [28, 24] assume the defender retains control over the fine-tuning process, we focus on settings where the defender cannot intervene after the weights are released or stolen. For a full review of current HFAs and threat models, we refer readers to [15].
@@ 7 Limitations
The primary limitation of RepNoise is that it is still possible to find ways to defeat it at higher learning rates and with more data (§ E.1). It is also sensitive to variations in hyperparameter choices (appendix J). We have evidence that RepNoise could be improved quite simply by doing more comprehensive hyperparameter searches and constructing larger defensive datasets. However, our method requires paired safe and unsafe examples which makes data collection more expensive and complex. Finally, while we did demonstrate comprehensive generalization across harmful subsets in question-answering tasks, we did not observe generalization from defences on harmful question-answering to attacks using toxic content generation (§ E.2)—as such, future work should focus on improving cross-domain defence.
While our empirical settings and attacks provide promising first directions for LLM immunization research, future work should invest in stronger attack settings to emulate worst-case attacks and investigate different types of harm. Finally, our work is limited to supervised fine-tuning attacks in LLMs. Additional settings in different modalities such as evaluating attempts at developing malicious agents through harmful reinforcement learning (e.g., “reverse DPO” [14]) are a critical topic for future research.
@@
We acknowledge the support of the Killam foundation, Digital Research Alliance of Canada, and the Vector institute. FR is supported by a Canada CIFAR Chair in AI.
@@ Appendix A Proofs and Mathematical Details
A.1 Deriving the approximate transition probability
@@
Motivated by the transfer learning question: "How can we predict the success of training a model originally trained on Task A to Task B", Achille et al. [18] present a transition probability that determines the likelihood an initial model with parameters $w_{0}$ can traverse the loss landscape to find the parameters $w_{f}$ that minimize the loss function $\mathcal{L_{D}}$ which represents a loss like causal language modelling loss on some dataset $D$.
They posit the following transition probability $p(w_{f},t_{f}|w_{0},t_{0})$ as equal to:
@@
Static potential measures how far an initial model is from the final configuration in terms of the difference between the loss of the model at the initial state $w_{0}$ and the final state $w_{f}$. This is due to the function $U(\cdot)$ which is the loss function given those parameters i.e. $\mathcal{L_{D}}(w)$. The constant $D$ comes from the author’s derivation of the original equation from a Wiener process. Minimizing this is where our Gradient Ascent loss comes from.
Reachability measures the likelihood of actually traversing the loss landscape. That is determined by integrating over the “difficulty” of reaching $w_{f}$ by integrating over all of the parameter configurations between $w_{0}$ and $w_{f}$ as well as the time steps it takes to reach $w_{f}$. “Difficulty” is measured by the terms $\dot{w}(t)^{2}+V(w(t))$. $\dot{w}$ is a stochastic differential equation that depends on the gradients of $\mathcal{L_{D}}$ with respect to the parameters $w$ with some stochastic function $\sqrt{2Dn(t)}$ i.e. $\dot{w}=\nabla\mathcal{L_{D}}(\theta)+\sqrt{2Dn(t)}$. This term simply expresses that the difficulty of a path is determined by how large the gradients are between parameter configurations. $V(w(t))$ is given by $\frac{1}{2}f(w_{t})^{2}+D\>\text{div}\>f(w_{t})$ where $f(\cdot)$ takes the gradients of the loss function with respect to the parameters $w$ at time step $t$. We also have an additional divergence term which measures properties of gradient flow on the path at $w_{t}$ reaching $w_{f}$.
@@
First, we observe that we can construct a simpler presentation for the main paper by ignoring the $D$ constants since estimating these properly would only make the transition probability smaller. Second, we make two strong assumptions. In the first strong assumption we replace the stochastic factor $\sqrt{2Dn(t)}$ with some constant which we set to 00 for simplicity. In practice, this stochastic factor will exist but when deriving our loss function we assume we won’t be able to control this factor so it isn’t important for deriving a transition probability that illustrates how to construct a loss function that minimizes it.
Our second strong assumption is that the divergence across all paths is set to some constant, in our case again it is replaced with 0. We realize that the divergence of the gradients of a loss function with respect to some parameters might play a major role in the transition probability: for example, if the divergence is always negative with a large magnitude, it is easy to rapidly traverse the loss landscape to $w_{f}$ and the transition probability would be much larger than otherwise expected without this term. Similarly, if the divergence is always positive with a large magnitude then the transition probability would be much smaller than expected. For our work, we only assert an approximate estimation of the transition probability where the divergence term and stochastic factors are necessary for a more accurate transition probability.
@@ A.2 Proof of Theorem 1
Let $\hat{Y}$ be the predicted label outputted by a representation model when feeding with input ${X}$, and suppose that ${Y}$ is a ground truth label for the input ${X}$. Denote the one-hot forms of $\hat{Y}$ and ${Y}$ as $\hat{\mathbf{Y}}$ and $\mathbf{Y}$, respectively. If the KL divergence loss $D_{\text{KL}}(\mathcal{P}(\hat{\mathbf{Y}})\|\mathcal{P}(\mathbf{Y}))$ increases, the mutual information between the representations $Z$ and ground truth outputs $Y$, $I(Z;Y)$ will decrease and vice versa.
First we point out that, in this context, the KL divergence loss over one-hot vectors is the same as the cross entropy loss (see the equivalence in § A.3). So we will refer to cross entropy loss increase as a way to decrease $I(Z;Y)$ and vice versa.
[…]
What about the static potential? It is easy to see (via the connection to increasing loss in theorem 2) that our newly derived minimizer also maximizes the static term of the transition probability since this minimizer also increases the cross entropy loss of the initial parameters. Therefore any minimizer of $I(Z;Y)$ is also a maximizer of the static potential further reducing the transition probability.
@@
A.3 Equivalence of KL and Cross Entropy
@@
$=1\cdot\frac{1}{Q(c)}+\sum_{i\neq c}0\cdot\>\text{log}\>\frac{0}{Q(i)}$ (11)
@@ B.1 Implementation of Representation Noising
13:     $\theta\leftarrow\theta-\eta\nabla_{\theta}\ell_{all}$
@@ B.2 Training details for the main results
For the results presented in § 4 and § 5, we use the following settings unless otherwise stated. We perform RepNoise on the same samples as the attack, while evaluations are always performed on unseen samples. We only use different samples for RepNoise in table 5 and table 17 to show its ability to generalise where we see that there is very little difference made whether we immunize on the same samples as the attack or not. For the BeaverTails attack, we train RepNoise for 1 epoch on 10k paired samples (10k harmful questions with harmful answers and 10k of the same questions with safe answers or refusals) using $\alpha=1$, $\beta=0.001$, learning rate (LR) $2\times 10^{-5}$. These settings were found after performing a grid search over $\alpha\in\{4,2,1,0.5,0.1\}$, $\beta\in\{2,1,0.1,0.01,0.001,0.0001\}$, $lr\in\{8\times 10^{-8},1\times 10^{-5},2\times 10^{-5},3\times 10^{-5},8\times 10^{-5},1\times 10^{-4},1\times 10^{-3}\}$, and 1, 2, and 4 epochs. The results of this search are presented in appendix J. For Decoding Trust RTP split, we used 4 epochs for immunization, a learning rate of $2\times 10^{-5}$, $\alpha=2$ and $\beta=4$. A random seed of 42 is used for all experiments unless otherwise stated. We use a cosine learning rate schedule with a $10\%$ warmup and use the Adam optimizer without any weight decay. Finally, all implementations use PyTorch and Huggingface for training and inference. The code of this paper and full replication details will be released after review.
@@ Appendix C Attack Setting Details
For the attack, we perform supervised fine-tuning to reduce causal language modelling loss on a harmful dataset (see section 2). We also evaluate the attacks from Qi et al. [8] and find that they do not increase the harmfulness of our base model appreciably ($0.04\rightarrow 0.06$ harmfulness on our classifier using all the same settings for the 100-shot attack) [^4] This could be due to differences in data; since the authors do not make their harmful samples available, we sample 100 harmful samples from BeaverTails.. Therefore we construct a set of stronger attacks. The strength of our attack depends on the number of samples and the learning rate $3\times 10^{-5}$, $6\times 10^{-5}$, $8\times 10^{-5}$}, which were chosen to allow for convergence while not degrading model quality (model disfluency occurs at $8\times 10^{-4}$). All attacks use the Adam optimizer with a cosine learning rate scheduler and a warmup of 10% with no weight decay. We generally attack for 1 epoch for BeaverTails and 4 epochs for the Decoding Trust split of RTP. As in Qi et al. [8] greedy sampling is used in all attacks. We illustrate the effects of sampling on our attacks in appendix J.

### v2

@@ 5 Mechanistic Analysis
Figure 5: Harmful probe accuracy on (a) base model and attacked model, (b) base model and models trained with RepNoise ($\beta=4$) and adversarial Loss, and (c) base model, RepNoise model and an attacked RepNoise model
@@ 6 Related Work
Preserving the effects of safety fine-tuning
Some prior work addresses the attenuation of safety fine-tuning’s influence on model behavior which typically occurs during benign fine-tuning. [78] achieve this for vLLMs with an instruction fine-tuning dataset which contains safety material and [49] do so for LLMs by modifying the prompt template used during fine-tuning. [32] use a modified LoRA algorithm for fine-tuning which maintains safety influence and [74] use a model fusion-based technique to get around the limitations of performing safety fine-tuning either before or after fine-tuning on tasks. Other solutions could benefit from methods that correct the general tendency for models to perform more poorly in some domains after being fine-tuned in others, such as the method presented in [51]. Though these methods may be sufficient for their stated goals, our work aims to prevent the effects of harmful fine-tuning, regardless of whether they come about from benign or harmful fine-tuning.
@@
Few works have attempted to defend against HFAs. Meta-learning approaches have been used to reduce the fine-tunability of Language Models for harmful classification tasks [28] and prevent image classification models from being fine-tuned on restricted domains [14]. However, meta-learning approaches can be uninterpretable and too computationally expensive for LLMs. [77] added a security vector during training to trick a model into thinking that it has already learned the harmful task. [36] keep embeddings close to the original embeddings by adding a perturbation loss called ‘Vaccination’ and [35] provides a similar defence by ensuring that weights don’t drift too far from original weights during training. While [77, 36, 35] assume the defender retains control over the fine-tuning process, we focus on settings where the defender cannot intervene after the weights are released or stolen. For a full review of current HFAs and threat models, we refer readers to [58].
@@ 7 Limitations
The primary limitation of RepNoise is that it is still possible to find ways to defeat it at higher learning rates and with more data (§ E.1). It is also sensitive to variations in hyperparameter choices (appendix J). We have evidence that RepNoise could be improved quite simply by doing more comprehensive hyperparameter searches and constructing larger defensive datasets. However, our method requires paired safe and unsafe examples which makes data collection more expensive and complex. Finally, while we did demonstrate across in-distribution harmful subsets in BeaverTails, we did not observe out-of-distribution generalization from defences on harmful question-answering to attacks using toxic content generation. Even smaller distribution shifts such as from defence using BeaverTails to an unseen harmful question-answering dataset HEX-PHI (§ E.2) can break RepNoise —as such, future work should focus on improving the generalization capabilities of RepNoise as it is unlikely that defenders will have access to samples with significant in-distribution overlap with attackers which limits the effectiveness of our proposed method.
While our empirical settings and attacks provide promising first directions for LLM immunization research, future work should invest in stronger attack settings to emulate worst-case attacks and investigate different types of harm. Finally, our work is limited to supervised fine-tuning attacks in LLMs. Additional settings in different modalities such as evaluating attempts at developing malicious agents through harmful reinforcement learning (e.g., “reverse DPO” [73]) are a critical topic for future research. We explored the implications of RepNoise for inference-time adversarial attacks (§ E.6) but future work should explore the robustness of RepNoise to additional types of attacks like latent adversarial attacks [11], activation engineering-based attacks [3] or adaptive attacks such as using decoding-time modifications [47] to circumvent our defence.
@@
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI) and Research Nova Scotia. Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada and the Vector Institute. FR is supported by a Canada CIFAR Chair in AI. We thank Ben Rank for providing the implementation of Vaccine that we use in this paper. The project on which this report is based was funded by the Federal Ministry of Education and Research under the funding code 16KIS2012. The responsibility for the content of this publication lies with the author
@@ Appendix A Proofs and Mathematical Details
A.1 Preliminaries
There are a number of terms used throughout the paper that require precise definitions where the main text is only able to provide a concise introduction. In this section, we will formally define the notions important to understanding the RepNoise algorithm and it’s theoretical analysis. First, we consider the weights of a deep neural network denoted $\theta$ as the parameters which are learned using some optimization procedure like that of the harmful fine-tuning procedure described in the main text in eq. 1, typically using stochastic gradient descent which we will return to shortly. We rely on the information bottleneck [64] view which states that neural networks form a Markov chain between the following random variables: the inputs of the network $X$, the representation of those inputs $Z$, and the outputs $Y$ which can predicted only from the representations $Z$. Here, we draw on the notion of representation from [2] which states that $Z$ is a representation of $X$ if in the chain described above $Z$ provides desirable properties for tasks involving $Y$. Here, the task of interest is predicting tokens $\hat{Y}$ such that those predicted tokens minimize some loss function $\mathcal{L}(\hat{Y},Y)$ over a reference token distribution $Y$. In this paper, the representations $Z$ will take the form of intermediate activations that are built up through linear transformations and activation functions in the transformer models that we use. Precisely, the neural network is the function $f_{\theta}(x)=\hat{y}$ parameterized by $\theta$ composed of activation functions $z_{i+1}=h(\theta_{i}\cdot z_{i})$ where $z_{i=0}$ is an initial embedding of $x$ such as through the use of a learned token embedding matrix common to LLM architectures and the final layer consists of an unembedding layer such as a weight matrix over a vocabulary space with a softmax function. In this process, since $\hat{y}$ is predicted through the intermediate $z_{i}$ activations then $z_{i}$ meets the criteria of being a representation of the initial input $x$. While the subspace that these activations span is completely determined by the weights $\theta$ themselves, we will not be referring to weights as representations in this paper.
To connect these representations back to an information-theoretic perspective, we will say that the information of any given representation $Z$ is the typical Shannon entropy measure of discrete random variables $H(Z)=\mathbb{E}[-\log p(Z)]$ where $p(x)=P(X=x)$. In this paper, we are concerned not with the information content of representations themselves but with the notion of mutual information: the amount of information one random variable gives us about another random variable. Formally, the (shannon) mutual information $I(X;Z)$ is given by the information content $H(X)$ minus the conditional entropy (or information content) $H(X|Z)$ for the formula $I(X;Z)=H(X)-H(X|Z)$. An equivalent distributional view of mutual information, which will will use later, can be given by the Kullback-Leibler divergence $I(X;Z)=\mathbb{E}_{x\sim P_{x}}[D_{KL}(P(z|x)||P(z)]$.
Given these tools we can represent a neural network as an information bottleneck which means that there is a Markov chain $Y\leftarrow Z\leftarrow X$ that has what is called a data processing inequality given by: $I(Y;X)\leq I(X;Z)$. We can also derive two ideal properties [2] of representations: sufficiency - the representations $Z$ have the same essential information in $X$ to pick out $Y$ i.e. $I(Z;Y)=I(X;Z)$; and minimality - the representations $Z$ should contain as little information of the input space as possible $\min I(X;Z)$. Finally, we present one more notion to formalize the idea of removing information from a representation. We say that a representation $Z$ has no information about a random variable $Y$ when the mutual information $I(Z;Y)$ is 0. The process of removing information more precisely means reducing the mutual information between these two random variables. As long as the sufficiency condition is met, removing information of representations $Z$ and outputs $Y$ doesn’t reduce the predictive capabilities of a neural network. In the context of unlearning and the preventing learning harmful text sequences, we want to remove mutual information between representations $Z$ and outputs $Y$ such that the sufficiency condition is not met and this is what we will mean precisely when we say removing information in the paper.
Finally, we need to present a few preliminaries from [1] in order to make the theoretical analysis clear. Below we will present a transition probability $p(s_{2},t_{2}|s_{1},t_{1})$ model which is a model of the likelihood of transitioning from one state $s_{1}$ to another $s_{2}$ at time step $t_{1}$ and $t_{2}$, readers familiar with reinforcement learning will recall that this is similar to the notion of a dynamics model of the environment. The transition probability model will consider the likelihood of transitioning between one set of model parameters $\theta_{1}$ to another set of parameters $\theta_{2}$ during the process of stochastic gradient descent over some loss function $L_{\theta_{i}}$. The loss landscape is the space ($\mathbf{R^{|}\theta|}$) of the value of a loss function $\mathcal{L}_{D}$ at every value of $\theta$. For simiplicity we are assuming this loss function is computed over all of the samples of a given dataset $D$ to construct our theoretical loss landscape object. We will develop a transition probability based on the static potential and reachability between two parameter sets. The difference $\mathcal{L}_{D\theta_{i}}-\mathcal{L}_{D\theta_{j}}$ between the loss for one parameter configuration $\theta_{i}$ and $\theta_{j}$ will be called the static potential below. Reachability will be developed precisely below. Paths or training trajectories in the loss landscape are the sequence of parameter configurations $\theta_{i}$ during a training procedure
A.2 Deriving the approximate transition probability
@@
Motivated by the transfer learning question: "How can we predict the success of training a model originally trained on Task A to Task B", Achille et al. [1] present a transition probability that determines the likelihood an initial model with parameters $\theta_{0}$ can traverse the loss landscape to find the parameters $\theta_{*}$ that minimize the loss function $\mathcal{L_{D}}$ which represents a loss like causal language modelling loss on some dataset $\mathcal{D}$. Here a Task is defined as a token output distribution over some dataset $\mathcal{D}$.
They posit the following transition probability $p(\theta_{*},t_{*}|\theta_{0},\theta_{0})$ as equal to:
@@
As mentioned in the preliminaries static potential $[U(\theta_{*})-U(\theta_{0})]$ measures how far an initial set of parameters is from a final set of parameters that minimize some loss term is given as the difference of loss between those two models $\mathcal{L}_{\mathcal{D}\theta_{*}}-\mathcal{L}_{\mathcal{D}\theta_{0}}$. The stochastic factor $D$ comes from the author’s derivation of the original equation from a Wiener process. Minimizing the static distance alone is where our Gradient Ascent loss comes from. Note that $\mathcal{D}$ refers to a dataset and $D$ refers to a stochastic factor.
Reachability measures the likelihood of traversing the loss landscape (as defined above). Reachability is determined by integrating over the “difficulty” of reaching $\theta_{*}$ by integrating over all of the parameter configurations between $\theta_{0}$ and $\theta_{*}$ as well as the time steps it takes to reach $\theta_{*}$ starting from each initial $\theta$ given by the outer integral. “Difficulty” is measured by the terms $\dot{w}(t)^{2}+V(w(t))$. $\dot{w}$ is a stochastic differential equation that depends on the gradients of $\mathcal{L_{D}}$ with respect to the parameters $\theta$ with some stochastic function $\sqrt{2n(t)}$ i.e. $\dot{w}=\nabla\mathcal{L_{\mathcal{D}}}(\theta)+\sqrt{2D(t)}$. This term simply expresses that the difficulty of a path is determined by how large the gradients are between parameter configurations. $V(w(t))$ is given by $\frac{1}{2}f(\theta_{t})^{2}+\nabla\cdot\>f(\theta_{t})$ where $f(\cdot)$ takes the gradients of the loss function with respect to the parameters $w$ at time step $t$. We also have an additional divergence term which measures properties of gradient flow on the path at $\theta_{t}$ reaching $\theta_{*}$.
@@
We observe that we can construct a simpler presentation for the main paper removing stochastic factors. If we properly estimate these factors this would only make the transition probability smaller due to the stochastic factor $D$ being the denominator of the reachability term as well as the stochastic factor $\sqrt{2D(t)}$ increasing the magnitude of the reachability term. In other words without stochastic factors, it is even more likely to reach $\theta_{*}$ and a minimizer of the deterministic transition probability would also minimize the stochastic transition probability.
Our second strong assumption is that the divergence across all paths is set to some constant, in our case again it is replaced with 0. We realize that the divergence of the gradients of a loss function with respect to some parameters might play a major role in the transition probability: for example, if the divergence is always negative with a large magnitude, it is easy to rapidly traverse the loss landscape to $\theta_{*}$ and the transition probability would be much larger than otherwise expected without this term. Similarly, if the divergence is always positive with a large magnitude then the transition probability would be much smaller than expected. For our work, we only assert an approximate estimation of the transition probability where the divergence term and stochastic factors are necessary for a more accurate transition probability.
@@ A.3 Proof of Theorem 1
Let $\hat{Y}$ be the predicted label output by an neural network with input ${X}$, and suppose that ${Y}$ is a ground truth next token label (in the form of a one-hot vector over a vocabulary) for the input ${X}$. If the KL divergence loss $D_{\text{KL}}(\mathcal{P}(\hat{Y})\|\mathcal{P}(Y))$ increases, the mutual information between the representations $Z$ and ground truth outputs $Y$, $I(Z;Y)$ will decrease.
First we point out that, in this context, the KL divergence loss over a one-hot target vector is the same as the cross entropy loss (see the equivalence in § A.4). So we will refer to cross entropy loss increase as a way to decrease $I(Z;Y)$.
Second, observe that we maximize cross entropy loss by taking the following gradient steps: $\theta_{t+1}=\theta_{t}+\eta\nabla\mathcal{L_{D}}\theta$. By theorem 2, increasing cross entropy loss increases the magnitude of the gradient of the loss function $\mathcal{L_{D}}$ with respect to the model parameters $\theta$. As established above, this magnitude is equivalent to the reachability term and therefore increasing cross entropy loss increases the reachability term which in turn minimizes the transition probability of finding the parameters $\theta$ that minimize $\mathcal{L_{D}}$.
By definition, maximizing cross entropy loss also increases the static potential term of the transition probability since it increases the loss of the initial parameters $\theta$ which will be used to attempt to train towards $\theta_{*}$.
The final step assumes the markov chain $Y\leftarrow Z\leftarrow X$ and the data processing inequality introduced in the preliminaries. We also assume that minmizing $I(Z;Y)$ also minimizes the cross entropy loss resulting in an increase in static potential and reachability, this can be seen from the definition of mutual information above. We must now to connect increase cross-entropy loss to a minmizer of $I(X;Z)$ and the minimization of $I(X;Z)$ to the minimization of both the static potential and reachability. From information bottleneck theory [71], we have the data processing inequality, $I(X;Y)$ will always be less than or equal to $I(X;Z)$ and therefore a minimizer of $I(X;Z)$ is a minimizer of $I(X;Y)$ which in turn is a minimize of $I(Z;Y)$. By theorem 2 we saw that increasing cross entropy (CE) loss decreases the quantity $I(Z;Y)$
@@
A.4 Equivalence of KL and Cross Entropy
@@
$=1\cdot\>\text{log}\>\frac{1}{Q(c)}+\sum_{i\neq c}0\cdot\>\text{log}\>\frac{0}{Q(i)}$ (11)
@@ B.1 Implementation of Representation Noising
13:     $\theta\leftarrow\theta-\eta\nabla_{\theta}\>\ell_{all}$
@@ B.2 Training details for the main results
For the results presented in § 4 and § 5, we use the following settings unless otherwise stated. For Table table 1 and table 2 we perform RepNoise and other defences on the same samples as the attack. This is the best-case defence scenario when the defender has access to the same samples as the attacker. For Table table 5 and table 20 we perform RepNoise on samples that are not used for the attack to show the ability to generalize where we see that there is very little difference made whether we immunize on the same samples as the attack or not illustrating that RepNoise still works when the defender does not have access to the same samples as the attacker but has access to the same domain. Post-attack evaluations are always performed on unseen samples.
For the BeaverTails attack, we train RepNoise for 1 epoch on 10k paired samples (10k harmful questions with harmful answers and 10k of the same questions with safe answers or refusals) using $\alpha=1$, $\beta=0.001$, learning rate (LR) $2\times 10^{-5}$. A batch size of $8$ is used throughout the experiments. These settings were found after performing a grid search over $\alpha\in\{4,2,1,0.5,0.1\}$, $\beta\in\{2,1,0.1,0.01,0.001,0.0001\}$, $lr\in\{8\times 10^{-8},1\times 10^{-5},2\times 10^{-5},3\times 10^{-5},8\times 10^{-5},1\times 10^{-4},1\times 10^{-3}\}$, and 1, 2, and 4 epochs. The results of this search are presented in appendix J. For Decoding Trust RTP split, we used 4 epochs for immunization, a learning rate of $2\times 10^{-5}$, $\alpha=2$ and $\beta=4$. A random seed of 42 is used for all experiments unless otherwise stated. We use a cosine learning rate schedule with a $10\%$ warmup and use the Adam optimizer without any weight decay. Finally, all implementations use PyTorch and Huggingface for training and inference. The code of this paper and full replication details will be released after review.
@@ Appendix C Attack Setting Details
For the attack, we perform supervised fine-tuning to reduce causal language modelling loss on a harmful dataset (see section 2). We also evaluate the attacks from Qi et al. [56] and find that they do not increase the harmfulness of our base model appreciably ($0.04\rightarrow 0.06$ harmfulness on our classifier using all the same settings for the 100-shot attack) [^4] This could be due to differences in data (we replicate the attack ”cross-domain” in § E.2; since the authors do not make their harmful samples available, we sample 100 harmful samples from BeaverTails.. Therefore we construct a set of stronger attacks. The strength of our attack depends on the number of samples and a chosen learning rate. For the main paper results we present a sampling of learning rates at the order of magnitude $10^{-5}$ since we observed that optimization using learning rates at lower than ($3\times 10^{-5}$) did not result in harmful models. Using learning rates at a higher order of magnitude ($5\times 10^{-4}$) often resulted in models with disfluent outputs. For the sake of convenience and concision, we arbitrarily select three learning rates ($3\times 10^{-5}$, $6\times 10^{-5}$, $8\times 10^{-5}$) but we present a more comprehensive analysis across learning rates in table 7. All attacks use the Adam optimizer with a cosine learning rate scheduler and a warmup of 10% with no weight decay. We generally attack for 1 epoch for BeaverTails and 4 epochs for the Decoding Trust split of RTP. As in Qi et al. [56] greedy sampling is used in all attacks. We illustrate the effects of sampling on our attacks in appendix J.
Model $1\times 10^{-5}$ $2\times 10^{-5}$ $4\times 10^{-5}$ $5\times 10^{-5}$
llama2-7b-chat 0.03 0.14 0.72 0.75
RepNoise 0.01 0.16 0.00* 0.04
$7\times 10^{-5}$ $9\times 10^{-5}$ $1\times 10^{-4}$ $5\times 10^{-4}$
llama2-7b-chat 0.77 0.74 0.74 0.00*
RepNoise 0.00* 0.08 0.08 0.00*
Table 7: An overview of attacks performed using learning rates ranging from ones too small to produce a viable attack to ones that are too large to maintain fluency of the models output generations. Asterisk indicates disfluency.